# A Statistical Framework of Watermarks for Large Language Models: Pivot, Detection Efficiency and Optimal Rules

### v1

@@ 5 Experiments
This section highlights our framework’s effectiveness through synthetic and real-data experiments, showcasing the practical utility of our proposed methods for detecting watermarks.
@@ 5.2 Real-World Examples
We present our numerical results with the OPT-1.3B models in Figure 7 (similar results for Sheared-LLaMA-2.7B are documented in Appendix D.2). Notably, both the average Type I error (leftmost column of Figure 7) and Type II error (rightmost column of Figure 7) curves are computed based on $500$ repeated experiments. Though the setup (mainly the way how each $\bm{P}_{t}$ is generated) is slightly different from our simulation studies, the conclusions are quite similar.
First, in most scenarios, all the detection methods maintain Type I errors within $0.04$ to $0.06$, aligning well with the expected $0.05$ level, particularly in texts over $100$ tokens long, as observed in our numerical experiments. Second, our methods, ${h_{\mathrm{gum},\Delta}^{\star}}(Y^{\mathrm{gum}})$ for Gumbel-max and ${h_{\mathrm{dif},\Delta}^{\star}}(Y^{\mathrm{dif}})$ for inverse transform, excel in minimizing Type II errors over other score functions. For Gumbel-max watermarks, the performance rankings follows our Theorem 3.3’s predictions: $h_{\mathrm{ars}}(Y^{\mathrm{gum}})$ records a lower Type II error, then $h_{\mathrm{log}}(Y^{\mathrm{gum}})$, with $h_{\mathrm{ind},\mathrm{e}^{-1}}(Y^{\mathrm{gum}})$ the highest type II error.
@@ 6 Discussion
Our work opens up several research directions in this emerging line of watermarks for LLMs. Perhaps the most immediate question is how to choose $\Delta$ for the distribution class ${\mathcal{P}}_{\Delta}$ in a manner adaptive to the LLM and specific problem instance. Another approach is to extend Proposition 2.1 to multiple distribution classes in recognition of the heterogeneity of NTP distributions across the token sequence. This can be done, for example, by adaptively clustering the empirical distributions of $\bm{P}_{t}$ into several clusters and using these to represent the distribution classes. Moreover, it would be a welcome advance to offer guidelines on how to choose pivotal statistics, which perhaps would allow the framework to maximize its utility since in this paper we basically fix the choice of pivotal statistics at the beginning. Finally, given that an unbiased watermark is probabilistically similar to sampling from a multinomial distribution, it would be of interest to consider other sampling schemes such as the alias method [63, 64] to develop new categories of watermarks.
@@ Acknowledgments
Appendix A Proof of Theorem 2.1
@@
By the strong law of large numbers, we know that
$\lim_{n\to\infty}\gamma_{n,\alpha}/n\leq\operatorname{\mathbb{E}}_{0}h(Y).$
@@
$\leq\exp(\gamma_{n,\alpha}\theta)\cdot\operatorname{\mathbb{E}}_{H_{1}}\exp\left(\sum_{t=1}^{n}-\theta h(Y_{t})\right)$
@@
By Lemma A.2, given that
we know that $R_{{\mathcal{P}}}(h)$ can also be written as
As a result, we can find a sequence of $\bm{P}_{k}\subset{\mathcal{P}}$ such that $\inf_{\theta\geq 0}Q(\theta,\bm{P}_{k})\to\inf_{\theta\geq 0}\sup_{\bm{P}\in\mathrm{cl}({\mathcal{P}})}Q(\theta,\bm{P})$ as $k\to\infty$; that is, $R_{\bm{P}_{k}}(h)\to R_{{\mathcal{P}}}(h)$ as $k\to\infty$.
$\lim_{n\to\infty}{\mathbb{P}}(T_{h}(Y_{1:n})=0|H_{1})^{1/n}=\mathrm{e}^{-R_{\bm{P}_{k}}(h)}\geq\mathrm{e}^{-(R_{{\mathcal{P}}}(h)+\varepsilon_{k}^{(1)})},$ (30)
where $\varepsilon_{k}^{(1)}\to 0$ as $k\to\infty$. It then follows that
${\mathbb{P}}(T_{h}(Y_{1:n})=0|H_{1})\geq\mathrm{e}^{-n(R_{{\mathcal{P}}}(h)+\varepsilon_{k}^{(1)}+\varepsilon_{n}^{(2)})},$
At the end of this section, we present the proof of Lemma A.2.
@@ Appendix B Proof for Gumbel-max Watermarks
B.1 Proof of Lemma 3.1
@@
$=\int_{0}^{r}{\mathrm{d}}r_{k}\int_{[0,1]^{|{{\mathcal{W}}}|-1}}{\bf 1}_{\{u_{j}\leq u_{{w}}^{P_{t,j}/P_{t,{w}}}~{}\text{for any}~{}j\neq{w}\}}{\mathrm{d}}u_{1}\ldots{\mathrm{d}}u_{{w}-1}{\mathrm{d}}u_{{w}+1}\ldots{\mathrm{d}}u_{|{{\mathcal{W}}}|}$
$=\int_{0}^{r}\prod_{j\neq k}u_{k}^{P_{t,j}/P_{t,{w}}}{\mathrm{d}}u_{w}$

### v2

@@ 5 Experiments
This section highlights our framework’s effectiveness through synthetic and real-data experiments, showcasing the practical utility of our proposed methods for detecting watermarks. [^17] The experiment code is collected in the GitHub repository https://github.com/lx10077/WatermarkFramework.
@@ 5.2 Real-World Examples
We present our numerical results with the OPT-1.3B models in Figure 7 (similar results for Sheared-LLaMA-2.7B are documented in Appendix D.2). Notably, both the average Type I error (leftmost column of Figure 7) and Type II error (rightmost column of Figure 7) curves are computed based on $500$ repeated experiments. Though the setup (mainly the way each $\bm{P}_{t}$ is generated) is slightly different from our simulation studies, the conclusions are quite similar.
First, in most scenarios, all the detection methods maintain Type I errors within $0.04$ to $0.06$, aligning well with the expected $0.05$ level, particularly in texts over $100$ tokens long, as observed in our numerical experiments. Second, our methods, ${h_{\mathrm{gum},\Delta}^{\star}}(Y^{\mathrm{gum}})$ for Gumbel-max and ${h_{\mathrm{dif},\Delta}^{\star}}(Y^{\mathrm{dif}})$ for inverse transform, excel in minimizing Type II errors over other score functions. For Gumbel-max watermarks, the performance rankings follow our Theorem 3.3’s predictions: $h_{\mathrm{ars}}(Y^{\mathrm{gum}})$ records a lower Type II error, then $h_{\mathrm{log}}(Y^{\mathrm{gum}})$, with $h_{\mathrm{ind},\mathrm{e}^{-1}}(Y^{\mathrm{gum}})$ the highest type II error.
@@ 6 Discussion
Our work opens up several research directions in this emerging line of statistical research on watermarks for LLMs. Perhaps the most immediate question is how to choose $\Delta$ for the distribution class ${\mathcal{P}}_{\Delta}$ when the knowledge of the underlying token distributions is limited. For instance, it would be valuable to explore methods for enhancing detection power when $\Delta$ is known in a distributional sense (see the discussion in Appendix D.4). Another approach is to extend Proposition 2.1 to multiple distribution classes in recognition of the heterogeneity of NTP distributions across the token sequence. This can be done, for example, by adaptively clustering the empirical distributions of $\bm{P}_{t}$ into several clusters and using these to represent the distribution classes. Moreover, it would be a welcome advance to offer guidelines on how to choose pivotal statistics, which perhaps would allow the framework to maximize its utility. In this paper, we fix the pivotal statistics from the beginning. However, we show that it is fundamentally difficult to identify optimal pivotal statistics that maximize our efficiency notion. The challenge lies in the nonconvex nature of the set of pivotal statistics, making the selection problem a nonconvex optimization problem that is currently beyond our reach (see Appendix A.3). Finally, given that an unbiased watermark is probabilistically similar to sampling from a multinomial distribution, it would be of interest to consider other sampling schemes such as the alias method [66, 67] to develop new categories of watermarks.
@@ Acknowledgments
Appendix A Proof Supporting the General Framework
A.1 Proof of Theorem 2.1
@@ Appendix A Proof Supporting the General Framework
By the strong law of large numbers, we know that as long as $\alpha\in(0,1)$,
$\limsup_{n\to\infty}\frac{\gamma_{n,\alpha}}{n}=\operatorname{\mathbb{E}}_{0}h(Y).$ (29)
We will prove this equation (29) at the end of the proof.
@@
$\leq\exp(\gamma_{n,\alpha}\theta)\cdot\operatorname{\mathbb{E}}_{H_{1}}\exp\left(\sum_{t=1}^{n}-\theta h(Y_{t})\right)$ (30)
@@
Proof of (29).
Under the null hypothesis, $Y_{t}$ are i.i.d. copies of $\mu_{0}$. By the strong law of large numbers, it follows that
$\frac{1}{n_{k}}\sum_{t=1}^{n_{k}}h(Y_{t})\to\operatorname{\mathbb{E}}_{0}h(Y)$
for any diverging subsequence $\{n_{k}\}_{k\geq 1}$ satisfying $\lim_{k\to\infty}n_{k}=\infty$.
By definition, $\gamma_{n,\alpha}$ is the theoretical critical value that makes the detection rule $T_{h}$ have a given Type I error of $\alpha$, i.e.,
${\mathbb{P}}_{H_{0}}\left(\sum_{t=1}^{n}h(Y_{t})\geq\gamma_{n,\alpha}\right)=\alpha.$
Assume, for contradiction, that $\limsup_{n\to\infty}\frac{\gamma_{n,\alpha}}{n}>\mathbb{E}_{0}[h(Y)]$. Then we could find a small positive number $\epsilon>0$ and a diverging subsequence $\{n_{k}\}_{k\geq 1}$ such that $\frac{\gamma_{n_{k},\alpha}}{n_{k}}\geq\mathbb{E}_{0}[h(Y)]+{\varepsilon}$ for sufficiently large $k$. This leads to a contradiction, as follows:
$\alpha$ $=\lim_{k\to\infty}{\mathbb{P}}_{H_{0}}\left(\sum_{t=1}^{n_{k}}h(Y_{t})\geq\gamma_{n_{k},\alpha}\right)$
$=\lim_{k\to\infty}{\mathbb{P}}_{H_{0}}\left(\frac{1}{n_{k}}\sum_{t=1}^{n_{k}}h(Y_{t})\geq\frac{\gamma_{n_{k},\alpha}}{n_{k}}\right)$
$\leq\lim_{k\to\infty}{\mathbb{P}}_{H_{0}}\left(\frac{1}{n_{k}}\sum_{t=1}^{n_{k}}h(Y_{t})\geq\operatorname{\mathbb{E}}_{0}h(Y)+{\varepsilon}\right)$
$={\mathbb{P}}_{H_{0}}\left(\operatorname{\mathbb{E}}_{0}h(Y)\geq\operatorname{\mathbb{E}}_{0}h(Y)+{\varepsilon}\right)=0.$
Similarly, if $\liminf_{n\to\infty}\frac{\gamma_{n,\alpha}}{n}<\mathbb{E}_{0}[h(Y)]$, we could find a small positive number $\epsilon>0$ and a diverging subsequence $\{n_{k}\}_{k\geq 1}$ such that $\frac{\gamma_{n_{k},\alpha}}{n_{k}}\leq\operatorname{\mathbb{E}}_{0}h(Y)-{\varepsilon}$ for sufficiently large $k$. This also leads to a contradiction:
$\alpha$ $=\lim_{k\to\infty}{\mathbb{P}}_{H_{0}}\left(\sum_{t=1}^{n_{k}}h(Y_{t})\geq\gamma_{n_{k},\alpha}\right)$
$=\lim_{k\to\infty}{\mathbb{P}}_{H_{0}}\left(\frac{1}{n_{k}}\sum_{t=1}^{n_{k}}h(Y_{t})\geq\frac{\gamma_{n_{k},\alpha}}{n_{k}}\right)$
$\geq\lim_{k\to\infty}{\mathbb{P}}_{H_{0}}\left(\frac{1}{n_{k}}\sum_{t=1}^{n_{k}}h(Y_{t})\geq\operatorname{\mathbb{E}}_{0}h(Y)-{\varepsilon}\right)$
$={\mathbb{P}}_{H_{0}}\left(\operatorname{\mathbb{E}}_{0}h(Y)\geq\operatorname{\mathbb{E}}_{0}h(Y)-{\varepsilon}\right)=1.$
@@
By Lemma A.2, given that
we know that $R_{{\mathcal{P}}}(h)$ can also be written as
As a result, we can find a sequence of $\bm{P}_{k}\subset{\mathcal{P}}$ such that
$\inf_{\theta\geq 0}Q(\theta,\bm{P}_{k})\to\inf_{\theta\geq 0}\sup_{\bm{P}\in\mathrm{cl}({\mathcal{P}})}Q(\theta,\bm{P})$
as $k\to\infty$; that is, $R_{\bm{P}_{k}}(h)\to R_{{\mathcal{P}}}(h)$ as $k\to\infty$.
$\lim_{n\to\infty}{\mathbb{P}}_{H_{1}}(T_{h}(Y_{1:n})=0)^{1/n}=\mathrm{e}^{-R_{\bm{P}_{k}}(h)}\geq\mathrm{e}^{-(R_{{\mathcal{P}}}(h)+\varepsilon_{k}^{(1)})},$ (31)
where $\varepsilon_{k}^{(1)}\to 0$ as $k\to\infty$. It then follows that
${\mathbb{P}}_{H_{1}}(T_{h}(Y_{1:n})=0)\geq\mathrm{e}^{-n(R_{{\mathcal{P}}}(h)+\varepsilon_{k}^{(1)}+\varepsilon_{n}^{(2)})},$
A.2 Proof of Proposition 2.1
In this subsection, we present the proof of Proposition 2.1 and establish the equation that $R_{{\mathrm{Simp}}({{\mathcal{W}}})}(h)=0$.
Proof of Proposition 2.1.
We start from the inequality (30). Since $\gamma$-fraction of $\bm{P}_{1},\ldots,\bm{P}_{n}$ belongs to ${\mathcal{P}}_{1}$ while the rest belongs to ${\mathcal{P}}_{2}$, it follows that
$1-\operatorname{\mathbb{E}}_{1}T_{h}(Y_{1:n})\leq\exp(\gamma_{n,\alpha}\theta)\cdot\exp(\gamma n\cdot\log\phi_{{\mathcal{P}}_{1},h}(\theta))+(1-\gamma)n\cdot\log\phi_{{\mathcal{P}}_{2},h}(\theta)).$
Hence,
$\limsup_{n\to\infty}\left[1-\operatorname{\mathbb{E}}_{1}T_{h}(Y_{1:n})\right]^{1/n}$ $\leq\inf_{\theta\in[0,\theta_{{\mathcal{P}},h})}\exp(\theta\operatorname{\mathbb{E}}_{0}h(Y)+\gamma\log\phi_{{\mathcal{P}}_{1},h}(\theta)+(1-\gamma)\log\phi_{{\mathcal{P}}_{2},h}(\theta))$
$\leq\exp(-[\gamma R_{{\mathcal{P}}_{1},h}+(1-\gamma)R_{{\mathcal{P}}_{2},h}])$
Lemma A.3.
If $\mu_{0}\in\{\mu_{1,P}:P\in{\mathrm{Simp}}({{\mathcal{W}}})\}$,
$R_{{\mathrm{Simp}}({{\mathcal{W}}})}(h)=0.$
Proof of Lemma A.3.
On one hand, we always have $R_{{\mathcal{P}}}(h)\geq 0$ for any ${\mathcal{P}}\subset{\mathrm{Simp}}({{\mathcal{W}}})$. The reason is given as follows. Recall that $R_{{\mathcal{P}}}(h)=-\inf\limits_{\theta\geq 0}f(\theta)$ where $f(\theta):=\theta\operatorname{\mathbb{E}}_{0}h(Y)+\log\phi_{{\mathcal{P}},h}(\theta)$. Clearly, $f(\theta)$ is convex in $\theta$ with $f(0)=0$. Hence, we always have
$R_{{\mathcal{P}},h}\geq-f(0)=0.$
On the other hand, we have $R_{{\mathrm{Simp}}({{\mathcal{W}}}),h}\leq 0$. The reason is given as follows. Note that $R_{{\mathcal{P}},h}$ is monotone in ${\mathcal{P}}$: if ${\mathcal{P}}_{1}\subset{\mathcal{P}}_{2}$, then $R_{{\mathcal{P}}_{2},h}\leq R_{{\mathcal{P}}_{1},h}$. If $\mu_{0}\in\{\mu_{1,P}:P\in{\mathrm{Simp}}({{\mathcal{W}}})\}$, without loss of generality, we assume $\mu_{0}=\mu_{1,\bm{P}_{0}}$. Since we must have $\{\bm{P}_{0}\}\subset{\mathrm{Simp}}({{\mathcal{W}}})$, it follows that
$R_{{\mathrm{Simp}}({{\mathcal{W}}}),h}\leq R_{\bm{P}_{0},h}=0.$
A.3 Difficulties of Finding the Optimal Pivot Statistic
At the end of this section, we illustrate that searching for the optimal pivotal statistic is not generally a convex problem. Overcoming these challenges would be significant and could be addressed in future research endeavors.
First and foremost, from an optimization perspective, the problem of finding the optimal pivotal statistic is in general non-convex. We will illustrate this non-convexity through concrete examples.
To set up the stage, we introduce some notation. For a given pivotal statistic $Y$, we first denote $J(Y)=\sup_{h}R_{{\mathcal{P}}}(h)$ to be the ${\mathcal{P}}$ efficiency for that pivotal statistic. We deduce
$J(Y)=\sup_{h}R_{{\mathcal{P}}}(h)$ $=-\inf_{h}\sup_{\bm{P}\in{\mathcal{P}}}\left(\operatorname{\mathbb{E}}_{0}h(Y)+\log\operatorname{\mathbb{E}}_{1,\bm{P}}\mathrm{e}^{-h(Y)}\right)$
$=-\sup_{\bm{P}\in{\mathcal{P}}}\inf_{h}\left(\theta\operatorname{\mathbb{E}}_{0}h(Y)+\log\operatorname{\mathbb{E}}_{1,\bm{P}}\mathrm{e}^{-h(Y)}\right)$
$=\inf_{\bm{P}\in{\mathcal{P}}}\sup_{h}\left(-\theta\operatorname{\mathbb{E}}_{0}h(Y)-\log\operatorname{\mathbb{E}}_{1,\bm{P}}\mathrm{e}^{-h(Y)}\right)$
$=\inf_{\bm{P}\in{\mathcal{P}}}D_{\mathrm{KL}}(\mu_{0}(Y)\|\mu_{1,\bm{P}}(Y)).$
The problem of finding the optimal pivotal statistic can be formally written down as:
${\rm maximize}~{}~{}J(Y)~{}~{}~{}\text{subject to}~{}~{}~{}\text{$Y=Y({w},\zeta)$ is a pivotal statistics}.$ (32)
Here we recall the notation that ${w}$ refers to the token, and $\zeta$ refers to the pseudorandom numbers. We give examples to demonstrate that (i) the functional $J$ is not convex nor concave, and (ii) the constraint set $Y$ being a pivotal statistic is not a convex set.
Consider the Gumbel-max watermark with two tokens, say, ${{\mathcal{W}}}=\{1,2\}$ (our examples extend to the general case $|{{\mathcal{W}}}|\geq 2$ by restricting the support of $\bm{P}$ to two). We denote the pseudorandom number $\zeta=(U_{1},U_{2})$ where $U_{1},U_{2}\sim{U}(0,1)$. Our examples take place in the simplest setting where ${\mathcal{P}}=\{\bm{P}\}$ is a singleton. In this setting, $J(Y)$ satisfies:
$J(Y)=D_{\mathrm{KL}}(\mu_{0}(Y)\|\mu_{1,\bm{P}}(Y)).$
Our first proposition states that $J(Y)$ is not necessarily convex or concave.
Proposition A.1 (The objective function is neither convex nor concave).
There is a singleton set $\{{\mathcal{P}}\}=\{\bm{P}\}$ with $Y_{1},Y_{2},Y_{3},Y_{4}$, each of which is a function of $({w},\zeta)$, such that
$J\left(\frac{Y_{1}+Y_{2}}{2}\right)<\frac{J(Y_{1})+J(Y_{2})}{2}~{}~{}\text{and}~{}~{}J\left(\frac{Y_{3}+Y_{4}}{2}\right)>\frac{J(Y_{3})+J(Y_{4})}{2}.$
Our second proposition demonstrates that the set of pivot statistics is not a convex set.
Proposition A.2 (The constraint is not convex).
There exist two pivotal statistics, $Y_{1}$ and $Y_{2}$, such that $\frac{Y_{1}+Y_{2}}{2}$ is not a pivotal statistic.
These two factors contribute to the significant challenges in solving the underlying optimization problem. We provide the proofs for the above two propositions in the following.
Proof of Proposition A.1.
We introduce the following functions and parameters:
$Y_{1}=U_{{w}},Y_{2}=U_{1},Y_{3}=U_{{w}}+U_{1},Y_{4}=|U_{{w}}-U_{1}|,~{}\text{and}~{}P_{1}=P_{2}=1/2.$
Table 1 presents some important calculation results which complete the proof. ∎
Function $Y$ $F_{0}(r)$ $F_{1}(r)$ $J(Y)=D_{\mathrm{KL}}(F_{0}\|F_{1})$
$Y_{1}=U_{{w}}$ $r$ $r^{2}$ $1-\log 2\approx 0.306$
$Y_{2}=U_{1}$ $r$ $r$ 0
$\frac{Y_{1}+Y_{2}}{2}=\frac{U_{{w}}+U_{1}}{2}$$\begin{cases}\frac{r}{2}+r^{2}~{}\text{if}~{}r\in[0,1/2]\\ \frac{5r-1}{2}-r^{2}~{}\text{if}~{}r\in[1/2,1]\end{cases}$$\begin{cases}3r~{}\text{if}~{}r\in[0,1/2]\\ 2-r~{}\text{if}~{}r\in[1/2,1]\end{cases}$ $\approx 0.0450295$
or $Y_{3}=U_{{w}}+U_{1}$
$Y_{4}=|U_{{w}}-U_{1}|$ $1-\frac{(1-r)^{2}}{2}$ $1-\frac{(1-r)^{2}}{2}$ 0
$\frac{Y_{3}+Y_{4}}{2}=\max\{U_{{w}},U_{1}\}$ $\frac{r+r^{2}}{2}$ $r^{2}$ $\approx 0.0996444$
Table 1: CDFs under different hypotheses and the efficiency rates of the considered functions. $F_{i}$ represents the CDF under hypothesis $H_{i}$. The domain for all the CDFs is $r\in[0,1]$. Here ${w}=1$ if and only if $\frac{\log U_{1}}{P_{1}}\geq\frac{\log U_{2}}{P_{2}}$. Recall that we use $P_{1}=P_{2}=1/2$ in the proof of Proposition A.1.
Proof of Proposition A.2.
Define $Y_{1}=U_{{w}}$ and $Y_{2}=U_{1}$. It is clear that $Y_{1}$ and $Y_{2}$ are pivotal statistics, as their distributions are independent of the distribution of ${w}$ under the null hypothesis $H_{0}$. In the following, we aim to compute the null CDF for $\frac{Y_{1}+Y_{2}}{2}$. Under $H_{0}$, we have $w\sim\bm{P}=(P_{1},P_{2})$ and $w\perp(U_{1},U_{2})$. As a result, it follows that for any $r\in[0,1]$,
${\mathbb{P}}_{H_{0}}\left(\frac{Y_{1}+Y_{2}}{2}\leq r\right)$ $={\mathbb{P}}_{H_{0}}\left(\frac{U_{w}+U_{1}}{2}\leq r\right)$
$=P_{1}\cdot{\mathbb{P}}(U_{1}\leq r)+P_{2}\cdot{\mathbb{P}}(U_{1}+U_{2}\leq 2r)$
$=P_{1}\cdot r+P_{2}\cdot\begin{cases}2r^{2}~{}\text{if}~{}r\in[0,1/2],\\ 1-2(1-r)^{2}~{}\text{if}~{}r\in[1/2,1].\end{cases}$
We conclude that $\frac{Y_{1}+Y_{2}}{2}$ is not a pivotal statistic because its null CDF depends on the distribution of ${w}$, i.e., $\bm{P}=(P_{1},P_{2})$. ∎
@@ Appendix B Proof for Gumbel-max Watermarks
B.1 Analysis for the pivotal statistic $Y^{\mathrm{gum}}$
@@
$=\int_{0}^{r}{\mathrm{d}}u_{{w}}\int_{[0,1]^{|{{\mathcal{W}}}|-1}}{\bf 1}_{\{u_{j}\leq u_{{w}}^{P_{t,j}/P_{t,{w}}}~{}\text{for any}~{}j\neq{w}\}}{\mathrm{d}}u_{1}\ldots{\mathrm{d}}u_{{w}-1}{\mathrm{d}}u_{{w}+1}\ldots{\mathrm{d}}u_{|{{\mathcal{W}}}|}$
$=\int_{0}^{r}\prod_{j\neq{w}}u_{{w}}^{P_{t,j}/P_{t,{w}}}{\mathrm{d}}u_{w}$
@@
At the end of this subsection, we argue that the pivotal statistic $Y^{\mathrm{gum}}_{t}=Y({w}_{t},\zeta_{t})=U_{t,{w}_{t}}$ in our manuscript is a good pivotal statistic. Our justification is that it approximates well the most optimistic likelihood ratio test statistic for an independent hypothesis test designed for watermark detection.
Suppose we have one generated token ($n=1$). Fix $t$ (and we drop the notation dependence on $t$). Consider the independence test regarding the token ${w}$ and the pseudorandom variable $\zeta$. Under $H_{0}$, ${w}$ and $\zeta$ are independent. Under $H_{1}$, ${w}$ and $\zeta$ are related by the decoder function ${\mathcal{S}}$. More precisely, we have the hypothesis testing problem:
$H_{0}:~{}{w}\perp\zeta~{}~{}\text{versus}~{}~{}H_{1}:~{}{w}={\mathcal{S}}(\bm{P},\zeta)~{}\text{for some $\bm{P}$}.$
This independent test is designed for watermark detection. We now compute the likelihood ratio test statistic. Let $\rho$ denote the probability density of $\zeta$.
  • Under $H_{0}$, the joint probability density of $({w},\zeta)$ is given by $P_{{w}}\cdot\rho(\zeta)$.
  • Under $H_{1}$, the joint probability density of $({w},\zeta)$ is given by ${\bf 1}_{{w}={\mathcal{S}}(\bm{P},\zeta)}\rho(\zeta)$.
The most optimistic log-likelihood ratio test statistic for this independent test is given by:
$\Lambda({w},\zeta)=\sup_{\bm{P}}\log\frac{{\bf 1}_{{w}={\mathcal{S}}(\bm{P},\zeta)}\rho(\zeta)}{P_{{w}}\cdot\rho(\zeta)}=\sup_{\bm{P}}\log\frac{{\bf 1}_{{w}={\mathcal{S}}(\bm{P},\zeta)}}{P_{{w}}}.$
We refer to $\Lambda({w},\zeta)$ as the most optimistic log-likelihood ratio test statistic because we take the supremum over all possible $\bm{P}$ to eliminate its dependence. Generally, we would reject $H_{0}$ if the log-likelihood ratio is sufficiently large. By using this most optimistic log-likelihood ratio, we adopt a more liberal approach when deciding to reject the null hypothesis.
We compute $\Lambda({w},\zeta)$ for the Gumbel-max watermark. Recall the notation $\zeta=(U_{w})_{{w}\in{{\mathcal{W}}}}$. We note that ${w}={\mathcal{S}}^{\mathrm{gum}}(\bm{P},\zeta)$ holds if and only if:
$\frac{\log\frac{1}{U_{w}}}{P_{{w}}}\leq\frac{\log\frac{1}{U_{k}}}{P_{k}}\quad\text{for any}\quad k\in{{\mathcal{W}}}.$
As a result, if ${w}={\mathcal{S}}^{\mathrm{gum}}(\bm{P},\zeta)$ holds, then:
$1=\sum_{k\in{{\mathcal{W}}}}P_{k}\leq\sum_{k\in{{\mathcal{W}}}}\frac{\log\frac{1}{U_{k}}}{\log\frac{1}{U_{w}}}P_{w}\implies P_{{w}}\geq\frac{\log\frac{1}{U_{w}}}{\sum_{k\in{{\mathcal{W}}}}\log\frac{1}{U_{k}}}.$
Using this relation, we deduce that
$\Lambda({w},\zeta)=\log\frac{\sum_{k\in{{\mathcal{W}}}}\log\frac{1}{U_{k}}}{\log\frac{1}{U_{{w}}}}.$
We argue that in the regime where $|{{\mathcal{W}}}|$ is large, this likelihood ratio statistic can be effectively approximated by
$\Lambda({w},\zeta)\approx\log\frac{|{{\mathcal{W}}}|}{\log\frac{1}{U_{w}}}$
as there is the law of large numbers where $\frac{1}{|{{\mathcal{W}}}|}\sum_{k\in{{\mathcal{W}}}}\log\frac{1}{U_{k}}\to 1$ for i.i.d. $U_{i}\sim U(0,1)$. Note, on the RHS, $\log\frac{|{{\mathcal{W}}}|}{\log\frac{1}{U_{w}}}$ is an increasing function of $U_{{w}}$, which is exactly the current pivotal statistic for the Gumbel-max watermark.
To conclude, we have shown that the pivotal statistic $Y^{\mathrm{gum}}_{t}$ in our manuscript, up to a monotone transform, approximates well the most optimistic likelihood ratio test statistic for the independent hypothesis test. This supports the effectiveness of $Y^{\mathrm{gum}}_{t}$ as a pivotal statistic for watermark detection. Thus, one contribution of our work is identifying the optimal transform $h$ based on $Y^{\mathrm{gum}}_{t}$ that achieves the highest possible statistical efficiency.